### v2

# HoneyBee: A Scalable Modular Framework for Creating Multimodal Oncology Datasets

Abstract
Developing accurate machine learning models for oncology requires large-scale, high-quality multimodal datasets. However, creating such datasets remains challenging due to the complexity and heterogeneity of medical data. To address this challenge, we introduce HoneyBee, a scalable modular framework for building multimodal oncology datasets that leverages foundation models to generate representative embeddings. HoneyBee integrates various data modalities, including clinical diagnostic and pathology imaging data, medical notes, reports, records, and molecular data. It employs data preprocessing techniques and foundation models to generate embeddings that capture the essential features and relationships within the raw medical data. The generated embeddings are stored in a structured format using Hugging Face datasets and PyTorch dataloaders for accessibility. Vector databases enable efficient querying and retrieval for machine learning applications. We demonstrate the effectiveness of HoneyBee through experiments assessing the quality and representativeness of these embeddings. The framework is designed to be extensible to other medical domains and aims to accelerate oncology research by providing high-quality, machine learning-ready datasets. HoneyBee is an ongoing open-source effort, and the code, datasets, and models are available at the project repository.
1 Introduction
The availability of large-scale public datasets has been a critical factor in advancing machine learning (ML) techniques, particularly in the development of foundation models [5, 13]. Datasets such as ImageNet [8], which contains over 14 million images across 20,000 categories, and COCO [20], which includes 330,000 images with over 200,000 labeled instances, have facilitated the training of deep learning models that achieve state-of-the-art (SOTA) performance on tasks such as object detection, semantic segmentation, and image captioning. In the context of oncology, multimodal medical datasets are likewise essential for developing ML models that can effectively capture the complex and heterogeneous nature of cancer. Medical data typically includes clinical records, imaging data (e.g., histopathology slides, radiology scans), molecular data (e.g., genomics, proteomics), and patient-reported outcomes. The integration of these diverse data modalities enables the identification of patterns and relationships that may not be apparent when analyzing individual modalities in isolation [35, 33, 4].
Multimodal datasets can support the development of ML models for various oncology applications, such as cancer screening, diagnosis, prognosis prediction, treatment response assessment, and post-treatment surveillance. Despite this need, there is a lack of large-scale, high-quality multimodal medical datasets for oncology research. The raw medical data collected from patients during studies are often distributed across multiple institutions and repositories, such as the Cancer Research Data Commons (CRDC) [15], Genomic Data Commons (GDC) [14], Proteomic Data Commons (PDC) [30], and Imaging Data Commons (IDC) [12], each with its own data formats, access protocols, and privacy constraints [32]. Integrating data from these disparate sources requires significant manual effort in terms of data harmonization, quality control, and metadata management. Furthermore, each data modality needs specialized preprocessing and feature engineering techniques to extract meaningful representations for ML tasks. For instance, histopathology slides may require tissue segmentation, color normalization, and tile extraction [21], while genomic data may involve quality control, normalization, and dimensionality reduction [34]. Additionally, clinical records often contain unstructured text data that needs to be processed using natural language processing (NLP) techniques [26]. These preprocessing steps are time-consuming, computationally expensive, and require domain-specific expertise. The scale and quality of existing publicly available multimodal oncology datasets vary considerably. Some datasets, such as The Cancer Genome Atlas (TCGA) [37], offer comprehensive multimodal data for multiple cancer types, including clinical records, imaging data, and molecular profiles, collected during large-scale studies. However, other datasets, like Med-MNIST [41], are more limited in scope, focusing on specific data modalities or providing smaller sample sizes, which may not be sufficient for developing robust ML models. Hence, there is a need for a framework that can efficiently aggregate and preprocess scattered raw public medical datasets to generate large-scale ML-ready feature representations that can be easily integrated into downstream ML pipelines.
To address the lack of critical large-scale multimodal medical datasets, we introduce HoneyBee, a modular and scalable framework for building ML-ready multimodal oncology datasets using open-source foundation models. The key objectives of HoneyBee are as follows:
  1. 1.
    Develop a set of standardized, modality-specific preprocessing pipelines for clinical records, imaging data, genomic information, and patient outcomes, ensuring consistency and reproducibility across datasets.
  2. 2.
    Capture complex patterns and relationships within and across data modalities by generating feature-rich embedding vectors from raw medical data using pre-trained foundation models.
  3. 3.
    Conduct a comprehensive evaluation of the HoneyBee framework using large-scale, oncology datasets by assessing performance in real-world applications.
The remainder of this paper is organized as follows. Section 2 reviews related work on multimodal oncology datasets, foundation models, and ML-ready datasets. Section 3 describes the HoneyBee framework in detail, including the data acquisition and integration process, embedding generation using the foundation models, data storage, and accessibility components. Section 4 presents the open and processed public datasets generated using the HoneyBee framework, focusing on TCGA [^1] The results shown here are in whole or part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga/. data source as a case study. Section 5 demonstrates the utility of HoneyBee datasets through use cases and experimental evaluations. Finally, we discuss the limitations, challenges, and future directions of the HoneyBee framework in Section 6.
2 Related Work
2.1 Challenges and Limitations of Current Public Datasets
Public datasets have been instrumental in advancing ML research, particularly in medical domains. However, these datasets, such as those integrated by the Multimodal Integration of Oncology Data System (MINDS) framework [32], often require extensive preprocessing to be ML-ready. MINDS consolidates data from various sources, such as TCGA [37] and TARGET [36], providing efficient query capabilities and facilitating interactive data exploration. Despite these capabilities, the datasets still face significant challenges, including inconsistent formats, missing data, and the need for domain-specific preprocessing techniques. Existing ML-ready datasets like Med-MNIST [41] and those available on Hugging Face Datasets [19, 7, 17, 23] have limited sample sizes and narrow focus on specific modalities. The need for preprocessing, harmonization, and normalization across different data sources remains a critical barrier to the effective use of these datasets in machine learning applications.
2.2 Foundation Models in Medical AI
Foundation models are pre-trained on extensive datasets and can be adapted to various medical tasks [27]. They integrate multiple data types, enhancing diagnostics, personalized treatment, and predictive analytics [43, 16, 44, 39, 25, 28, 31]. Models like UNI [6] and REMEDIS [3] have shown effectiveness in pathology and other medical fields. However, the development of truly multimodal foundation models that utilize raw medical data from various sources, such as whole slide images (WSI), molecular sequences, radiology scans, and electronic health records (EHR), is limited [2]. This limitation is primarily due to the lack of comprehensive multimodal datasets that integrate these diverse data types.
3 The HoneyBee Framework
The HoneyBee framework consists of three main components: data acquisition and integration, embedding generation, and data storage and accessibility, as illustrated in Figure 1. We discuss these components in detail below.
Figure 1: The HoneyBee workflow starts with accessing the public datasets, including data aggregation, cataloging, and metadata tagging using scripts and libraries from MINDS. The curated cohort dataset is then processed through modality-specific pipelines and pre-trained models in the HoneyBee framework. The resulting embeddings, as well as the raw data, are made publicly accessible to support the development of ML applications.
3.1 Data Acquisition and Integration
HoneyBee extends the data integration capabilities of the MINDS framework [32] by incorporating additional preprocessing steps to ensure data quality and compatibility. The data in MINDS is collected from various sources, including public repositories, clinical institutions, and research collaborations. Key data modalities incorporated in HoneyBee datasets include:
Text: Text data includes structured and unstructured medical reports, such as pathology reports, radiology reports, and clinical notes, generally stored in electronic health records (EHRs). These datasets provide valuable information about patient diagnoses, treatments, and outcomes. HoneyBee employs NLP techniques to extract relevant features and normalize the text data for downstream ML. The collected data undergoes integration steps to harmonize heterogeneous datasets and align data elements across modalities. Metadata alignment consolidates metadata from various text sources to ensure consistency by mapping different terminologies to a common ontology and standardizing data types. Data cleaning handles missing values, outliers, and inconsistencies, which may be imputed using statistical methods or domain-specific rules. Harmonization adheres to standards like Fast Healthcare Interoperability Resources (FHIR) [1] for EHRs to ensure interoperability and facilitate data exchange, including standardizing data formats and protocols to convert data from different formats to a common standard.
Imaging: Imaging data may include (1) digitized histopathology slides WSIs, and (2) radiology scans, performed using computed tomography (CT), magnetic resonance imaging (MRI), positron emission tomography (PET), Ultrasound, X-ray, etc. These imaging modalities capture detailed visual information about tumor morphology, size, location, and progression. HoneyBee utilizes SOTA computer vision models to process and analyze these high-dimensional imaging data. The collected data undergoes integration steps to harmonize heterogeneous datasets and align data elements across modalities. Metadata alignment ensures consistency in metadata by mapping terminologies to a common ontology and standardizing data types. Data cleaning addresses missing values, outliers, and inconsistencies through statistical imputation or domain-specific rules. Normalization standardizes pixel values across different imaging sources using data collecting protocols and recording information, available in the metadata of radiology data storage file format, Digital Imaging and Communications in Medicine (DICOM). Harmonization utilizes DICOM standards to ensure comparability and interoperability across imaging datasets, including standardizing data formats and protocols. For WSIs, additional preprocessing steps, such as stain normalization, patching, and background removal, are performed to ensure data quality and compatibility.
Molecular: Molecular data types provide insights into the underlying biological mechanisms of cancer at the genomic, transcriptomic, and proteomic levels. HoneyBee incorporates specialized bioinformatics pipelines to preprocess and integrate these complex molecular data. The collected data undergoes integration steps to harmonize heterogeneous datasets and align data elements across modalities. Metadata alignment consolidates different terminologies into a common ontology and standardizes data types. Data cleaning handles missing values, outliers, and inconsistencies through statistical imputation or domain-specific rules. Normalization employs methods to standardize numerical features, ensuring consistent data representation. Harmonization ensures interoperability and data exchange by adhering to established standards and addressing inconsistencies within and between datasets.
3.2 Embedding Generation
HoneyBee utilizes foundation models to generate embeddings from raw medical data. The embedding generation process involves several key steps:
3.2.1 Selection of Foundation Models
The HoneyBee framework includes foundation models tailored for different medical modalities:
  • For radiology scans (CT, MRI, PET), the REMEDIS model, trained on large-scale medical imaging datasets, generates embeddings that capture spatial features and structural information.
  • For WSIs from digitized histopathology specimens, the TissueDetector model detects and segments regions of interest, while the UNI model, a ViT-based encoder pre-trained on the Mass-100K dataset, generates embeddings that capture visual features and patterns [6].
  • For textual data (medical reports, clinical notes), language models from the Hugging Face library are used.
  • For molecular data (gene expression, DNA methylation, protein expression, mutations, miRNA), the SeNMo model, a deep learning framework, integrates various data types and generates unified latent embeddings [34].
3.2.2 Preprocessing of Raw Medical Data
Preprocessing in HoneyBee ensures the compatibility of raw imaging, textual, and molecular data with foundation models. Processing images involves resizing images to fit the input requirements of the models while preserving the aspect ratio and minimizing information loss, typically using adaptive techniques. Normalization scales pixel values to a standard range, ensuring uniformity across datasets. Data augmentation techniques such as rotation, flipping, and cropping are applied to enhance the diversity of the dataset, improving model robustness. Tissue detection models are employed to identify and extract regions of interest from WSIs by reading regions at various resolutions and processing them using ML models designed for tissue detection. For textual data, the preparation begins with tokenization, converting text into a sequence of tokens suitable for model processing. This process addresses complex medical terminology to ensure effective representation learning. Optical Character Recognition (OCR) is used to extract text from images of documents, including scanned pathology reports. The OCR process includes steps like grayscaling, resizing, and thresholding to enhance accuracy. The extracted text undergoes cleaning to remove noise and irrelevant characters, including extra whitespace, newline characters, and non-alphanumeric symbols. For long documents, the text is split into manageable chunks using techniques that ensure proper handling of chunk size and overlap. For molecular data, the preparation involves several steps to prepare the data for analysis. First, missing values are addressed using robust imputation methods to ensure completeness. Features with no variance are removed, as they do not contribute to the model’s discriminative capabilities. Redundant data points are identified and eliminated through deduplication algorithms to maintain data integrity. Highly correlated features are dropped to avoid multicollinearity, which could negatively impact model stability and interpretability. Missing data points are then imputed using advanced techniques, such as matrix factorization and deep learning-based methods, ensuring that the imputed values are consistent with the underlying data distribution.
3.2.3 Generation of Embeddings
Each preprocessed data sample is passed through the foundation model, which produces a fixed-length embedding vector. For example, the embedding dimensions can range from 48 to 2048, depending on the model architecture. HoneyBee utilizes GPU acceleration and distributed computing, when available, to efficiently generate embeddings for large-scale datasets. The generated embeddings, along with associated metadata, are stored in a structured format to facilitate downstream tasks such as similarity search, clustering, and ML model training. HoneyBee employs efficient data compression and indexing techniques to optimize the storage and retrieval of high-dimensional embeddings.
Figure 2: Pathology Workflow: WSIs are processed into $512\times 512512 × 512$ pixel tiles, followed by tissue detection and embedding generation. Radiology Imaging Workflow: CT, MRI, and PET scans undergo normalization, standardization, slicing, and embedding generation. Molecular Data Workflow: DNA methylation, gene expression, protein expression, DNA mutation, and miRNA expression data are preprocessed (removing NaNs and low-expression genes), and embeddings are generated. Clinical Data Workflow: Text from EHRs and reports is extracted, processed, and embeddings are generated. All processed data is integrated into a multimodal embedding dataset and optionally stored in vector databases.
3.3 Data Storage and Accessibility
The generated embeddings and tabular data are stored using the Hugging Face datasets library, which provides a standardized interface for data access and processing [19]. The datasets are organized into a structured format, containing the embeddings, metadata, and labels (if available). PyTorch DataLoaders are employed to efficiently load and iterate over the datasets during model training and evaluation, handling tasks such as batching, shuffling, and parallel processing. Additionally, HoneyBee datasets can be integrated into vector databases such as Faiss and Annoy [11, 38] to enable fast similarity search, nearest neighbor retrieval, and clustering on the high-dimensional embedding vectors. These databases are optimized for efficient querying and retrieval of embeddings based on similarity metrics such as Euclidean distance or cosine similarity [40]. By deploying vector databases, researchers can quickly identify similar samples, perform data exploration, and retrieve relevant subsets based on embedding similarity, facilitating various downstream tasks such as retrieval augmented generation (RAG) [42]. The structured storage and accessibility components of HoneyBee ensure that the generated embeddings and associated data are readily available for researchers and practitioners to use in their ML pipelines and downstream applications.
4 Datasets
[…]
Table 1: Overview of models used for different data modalities in HoneyBee. The table lists the types of data (Modality), the corresponding machine learning models utilized (Model), the shape of the embeddings generated by these models (Embedding Shape), and the number of samples for each modality (Sample Count).
Modality Model Embedding Shape Sample Count
Clinical Data GatorTron-medium [43] [1024] 11,428
Pathology Reports GatorTron-medium [43] [1024] 11,208
Pathology Images UNI [6] [patches, 1024] 30,075
Radiology Images REMEDIS [3] [patches, 7, 7, 2048] 11,869
Molecular Data SeNMo [34] [48] 13,804
We preprocess the raw data for each modality, feed it into the selected model at inference time, and generate a fixed-length embedding vector per sample, per modality. The embeddings are stored with associated metadata using the Hugging Face datasets library. To facilitate easy access and utilization of the generated embeddings, we have made the processed TCGA dataset, comprising patient embeddings, publicly available on the Hugging Face platform (https://huggingface.co/datasets/Lab-Rasool/TCGA) under an open Creative Commons Attribution Non Commercial No Derivatives 4.0 license.
5 Use Cases
To validate the effectiveness of the HoneyBee framework in generating meaningful multimodal datasets, we conducted a series of experiments assessing the quality and utility of the embeddings on a downstream ML task. We used all the clinical text data generated from the 33 cancer sites in the TCGA dataset and extracted embeddings using the GatorTron (gatortron-medium) and BERT (bert-base-uncased) models [43, 10]. We trained a random forest classifier to classify the cancer type using the embeddings available in the HoneyBee Hugging Face repository. In addition to showcasing the capability of generating embeddings directly from the models, we demonstrate how parameter-efficient fine-tuning (PEFT), which has gained popularity recently with the explosion of large language model research [18, 9], can make the embedding models more suited for the task. The experiments were conducted using an Nvidia RTX 3090 GPU with 24GB VRAM, 32GB RAM, and a Ryzen 5950X 16-core CPU.
5.1 Analysis of the Extracted Embedding in HoneyBee Framework
We began by analyzing the quality of the extracted embeddings and visualizing them using t-SNE. This dimensionality reduction technique maps high-dimensional data to a lower-dimensional space while preserving local structure. Figure 3 show the t-SNE plots of the embeddings generated by the pre-trained and fine-tuned versions of the GatorTron (https://huggingface.co/Lab-Rasool/gatortron-base-tcga) and BERT (https://huggingface.co/Lab-Rasool/bert-base-uncased-tcga) models, respectively. Each point represents a patient’s clinical information, while the colors correspond to the different TCGA project IDs. In both cases, the fine-tuned embeddings exhibit more visual separation between the different project IDs compared to the pre-trained embeddings, suggesting that the fine-tuning process has successfully adapted the language models to capture the nuances and distinguishing features of the various cancer types in the TCGA dataset. We employed the adapter-based fine-tuning approach to fine-tune the models to showcase the HoneyBee framework’s capability for PEFT, where a small set of low-rank adapters are inserted into the transformer (query, key, and value) linear layers of the pre-trained model, and only these adapter layers are updated during fine-tuning. In contrast, the pre-trained model parameters remain frozen.
Figure 3: t-SNE visualization of pre-trained and fine-tuned embeddings generated by the GatorTron and BERT models. Each subplot represents the embeddings for a specific model configuration: (a) t-SNE of GatorTron Pre-trained Embeddings, (b) t-SNE of GatorTron Fine-tuned Embeddings, (c) t-SNE of BERT Pre-trained Embeddings, and (d) t-SNE of BERT Fine-tuned Embeddings. Each point corresponds to a clinical note, with colors indicating different TCGA studies. Fine-tuned embeddings demonstrate improved separation between studies, indicating enhanced model performance in distinguishing between cancer types post-fine-tuning.
5.2 Downstream ML Model Training
To evaluate the effectiveness of the generated embeddings for cancer-type classification, a random forest classifier was trained. The clinical dataset from TCGA, including clinical text data and project IDs, was loaded and converted into a pandas DataFrame. Two models, GatorTron-medium and BERT-base-uncased, were used for generating embeddings, with both pre-trained and fine-tuned versions created. The clinical text data was tokenized and passed through the models to generate fixed-length embeddings. The embeddings and corresponding project IDs were split into training (80%), and test (20%) sets and a random forest classifier with 100 estimators was trained on the training set embeddings. The trained classifier was evaluated on the test set embeddings, and accuracies were calculated. To ensure robustness and account for variability, the experiment was run multiple times with different random seeds. Specifically, we conducted 10 runs with different random seeds for both the pre-trained and fine-tuned models. The mean accuracy and standard deviation of these runs were calculated to provide error bars for the performance metrics. Table 2 shows the classification accuracies achieved by the random forest classifier using both pre-trained and fine-tuned embeddings from the GatorTron and BERT models. The fine-tuned models exhibited improved performance over the pre-trained models, with the GatorTron model demonstrating superior accuracy in both scenarios.
Table 2: Classification and retrieval accuracies for different cancer types using pre-trained and fine-tuned embeddings. The GatorTron-medium model outperforms the BERT-base-uncased model in both pre-trained and fine-tuned settings, with significant improvements observed after fine-tuning.
Model Pre-trained Fine-tuned
Accuracy Retrieval Accuracy Retrieval
[…]
5.3 Retrieval Benchmark
In addition to classification, we conducted a retrieval benchmark to assess the ability of the embeddings to capture similarity between clinical texts from the same cancer type. This benchmark used FAISS [11] for similarity search and evaluating the retrieval of patient records with matching project IDs. We ran 100 trials where a random patient’s clinical embedding was selected as a query, and the nearest neighbors were retrieved from the embedding space. The percentage of correct matches (i.e., retrieved patients having the same project ID as the query) was calculated. This process was repeated 100 times, and the mean and standard deviation of the retrieval accuracy were computed. The results, presented in Table 2, show significant improvements in retrieval accuracy for fine-tuned models compared to pre-trained models, with the GatorTron [43] model again outperforming BERT in both settings. These retrieval benchmarks demonstrate the effectiveness of the HoneyBee framework in generating embeddings that support both classification and retrieval tasks in medical research and clinical data analysis.
6 Discussion and Conclusion
The HoneyBee framework integrates multimodal data and uses representation learning techniques to create ML-ready datasets for oncology research. The use cases demonstrate the effectiveness of the generated embeddings in capturing information from raw medical data and their utility in tasks such as cancer-type classification. The t-SNE visualizations (Figure 3) show that fine-tuned embeddings separate different cancer types better than raw embeddings, suggesting that fine-tuning adapts the language models to capture distinguishing features of cancer types in the TCGA dataset. Comparing the performance of the GatorTron model to BERT highlights the importance of selecting appropriate foundation models for each data modality. The HoneyBee framework allows for the incorporation of various data modalities and the flexibility to extend to other disease areas. It provides a standardized pipeline for data integration, pre-processing, normalization, harmonization, embedding generation, and storage, facilitating the creation of datasets that can accelerate model development in oncology and other medical domains. This work addresses the need for large-scale, high-quality, ML-ready datasets in the oncology domain, which can drive the development of innovative models and analyses. However, the interpretability and trustworthiness of the generated embeddings require further investigation to facilitate adoption in clinical settings [29]. The TCGA dataset may contain biases due to the patient selection criteria and data collection processes, which could impact the generalizability of the models. HoneyBee invites collaborators to contribute to the ongoing open-source effort.